THE ANALYSIS OF HIGH RESOLUTION SPECTRA OF ASYMMETRICALLY DEUTERATED METHOXY RADICALS CH_{2}DO AND CHD_{2}O

MING-WEI CHEN, JINJUN LIU,^{a} \underline{DMITRY G. MELNIK}, and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210; ROBERT F. CURL, Department of Chemistry and Rice Quantum Institute, Rice University, Houston TX, 77005; C. BRADLEY MOORE, Department of Chemistry, University of California, Berkeley CA, 94720.

Asymmetric deuteration of methoxy radical reduces the symmetry of normal vibrations, while the electronic wavefunction essentially retains its threefold symmetry. Although the vibrational modes are no longer degenerate, the asymmetry induced by deuteration can be treated as a perturbation and the molecule can still be analyzed using theory similar to that which has been employed for the isotopically symmetric CH_{3}O and CD_{3}O. The treatment of the rotationallyresolved spectra of these molecules is especially interesting since the angular momentum is no longer directed along one of the principal inertial axes. We have recently proposed the methods^{b} for handling this kind of a problem. This presentation describes the application of these methods to the global analysis of the high resolution data on CH_{2}DO and CHD_{2}O radicals, including microwave spectra^{c}, laser induced fluorescence (LIF) and stimulated emission pumping (SEP) spectra.^{d} The analysis of these data along with those existing for the symmetric species^{e} provides quantitative insight into the mechanism of Jahn-Teller interaction, spin-orbit interaction, and the effects of deuteration and symmetry reduction on these phenomena.

apresent address: Laboratory of Physical Chemistry, ETH Zurich, Wolfgang-Pauli-Str. 10, 8093 Zurich, Switzerland

bD.Melnik, J.Liu, R.F.Curl, and T.A.Miller, Mol.Physics, 105, 529 (2007)

cD. Melnik et al. 59^{th} International Symposium on Molecular Spectroscopy, WJ09, (2004)

dJ.Liu et al. 62^{nd} International Symposium on Molecular Spectroscopy, RF02, (2007)

esee preceding talk.